### v3

@@ 2 AI for Network
The most fundamental goal of the integration of wireless communication and AI lies in enhancing the performance, efficiency of the network itself and the service experience of users by virtue of AI, which is known as AI4NET [36]. At this stage, the focus is on how to utilize AI algorithms to optimize communication algorithms and network functions. For example, AI can be used to optimize the signal modulation and demodulation process, making signal transmission more accurate and efficient; with the help of AI, network resources can be intelligently allocated to achieve load balancing; and an automated O&M management mode can be constructed to improve O&M efficiency and reduce costs. Under this mode, the introduction of AI will not have a significant impact on the original network architecture. It mainly applies and improves specific identified communication problems by training AI algorithm models, just like precisely implanting intelligent patches into the network to gradually improve the intelligence level of the network while maintaining the stability of the architecture.
The conceptual diagram of AI4NET is presented in Figure 5, where the wireless AI possesses capabilities such as feature extraction, prediction, adaptation, optimization, real-time processing, correlation, and scene clustering. These capabilities form the foundation for AI-enabled wireless networks and can be directly deployed on the BS side and the CN to empower wireless communication. In the user environment, AI optimizes conventional methods to enhance network performance on various terminal devices. On the network side, AI takes advantage of its capabilities to improve the quality of end-to-end services, optimize the functions of network elements, conduct automated management of the upper-layer network, and boost network O&M efficiency. Meanwhile, the advent of AI has also given rise to a series of new services aimed at improving and serving users in a better manner.
AI4NET has initiated relevant research and applications in 5G. In 6G, with the more mature AI technologies represented by DL and the emergence of new infrastructures integrating connectivity and computing power, the relevant applications will become more abundant and sophisticated and may further evolve in depth. Next, this section will enumerate some cases from two aspects, namely air interface performance enhancement and network O&M improvement, to better illustrate the concept of AI4NET.
@@ 2.1 Air Interface Performance Enhancement
The performance of the wireless air interface will directly affect the user’s communication experience. Finding a better air interface algorithm to optimize air interface performance has always been the goal of industry insiders. However, user demands are increasing with the continuous increase of communication services. conventional algorithms incur complexity through complex matrix operations and reach the performance bottleneck.
The booming development of AI has provided a new direction for air interface transmission in communication. AI can fit complex and nonlinear mapping relationships. Compared with conventional methods, AI is more suitable for applications in complex time-varying wireless environments. Thus, AI can replace conventional communication methods to improve wireless network performance significantly. In this section, we show how AI technology can improve the transmission performance in CSI feedback, orthogonal frequency division multiplexing (OFDM) receivers, beam management, and wireless localization.
@@ 2.1.1 AI-Based CSI Feedback Algorithm
Within the framework of 6G communication technology, the feedback overhead experiences a sharp increase along with the significant rise in the number of antennas. Due to the extremely high demand of 6G networks for refined resource utilization, the uplink resources become increasingly scarce and precious. Consequently, ensuring the accuracy of channel reconstruction and minimizing the CSI feedback cost is one of the bottleneck problems that need to be surmounted. The utilization of AI technology in CSI feedback is capable of profoundly dissecting these elaborate channel characteristics. Through training with a substantial quantity of labeled samples, the AI model can acquire the characteristic patterns of the channel within diverse fading environments, interference scenarios, and multipath circumstances. In contrast to conventional CSI estimation approaches relying on statistical theory or simple linear models, AI-based methods exhibit stronger generalization competencies and can estimate CSI with greater precision in complex real-world communication settings, effectively diminishing the estimation error. This highly accurate CSI estimation facilitates the communication system to fine-tune transmission parameters, such as modulation and coding schemes and transmission power, more accurately, thus augmenting the spectral efficiency and transmission reliability of the system.
The massive CSI data and the inherent random characteristics of CSI make it preferable to introduce AI for designing a new CSI feedback mechanism. For instance, the nonlinear characteristics of DL can be utilized to efficiently extract the features of CSI, and the original CSI data can be transformed into a more compact feature representation, thereby reducing the communication overhead in the CSI feedback process. An example of its implementation is the autoencoder architecture, which employs neural networks to extract and compress the features of CSI channel information. The authors of [37] were the first to apply DL to CSI feedback first. The proposed CsiNet scheme utilizes convolution to extract channel features, compress and reconstruct channels, and it is superior to the CSI feedback scheme based on compressed sensing in terms of feedback accuracy and computational complexity. The encoder of CsiNet compiles the high-dimensional CSI into code words, while the decoder is responsible for decoding the code words back into the original CSI. The internal structures of encoders and decoders of different algorithms vary, resulting in differences in channel reconstruction performance. As shown in Figure 6, Huawei Technologies Co., Ltd. also demonstrated the huge advantages of AI-based CSI feedback over conventional codebook approach through system-level simulations and AI with additional information can effectively improve the performance on the square of the generalized cosine similarity (SGCS) and throughput [38].
@@
Figure 6: The performance of R16 Type II codebook, AI-SF and AI-SFT on SGCS and throughput gain. AI-SF means the AI/ML-based spatial-frequency compression method, which transformer is used as the backbone of both encoder and decoder. AI-SFT is AI-SF with additional past CSI.
The studies in [39] increased the size of the convolutional kernel based on CsiNet to improve the perceptual field of view of the convolutional layer, which was conducive to extracting the sparsity of the channel and further enhancing the accuracy of channel reconstruction. Although a larger perceptive field of view could effectively extract the sparsity of CSI, a smaller convolution kernel could extract finer features from CSI. A DL-based CSI compression and quantization method was developed in [40] considering high compression ratios. A multi-rate CSI compression framework was designed in [41] to improve the generalization of the model in the field of transfer learning. In [42], a non-local neural network was introduced based on the CsiNet network to capture a wide range of dependencies, and the accuracy of channel recovery was improved compared to CsiNet. The CRNet was proposed in [43], which used convolution kernels of different sizes for channel feature extraction and recovery, reducing the computation and improving the accuracy of channel reconstruction.
The above-mentioned literature aimed at using DL to improve the compression and recovery accuracy of CSI. However, in order to better deploy the AI-based CSI feedback model in the actual wireless environment, model monitoring, model updating and AI-based signaling process management are hot issues that require our attention in future studies.
@@ 2.1.2 AI-Based OFDM Receiver
The wireless communication transmitter and receiver are divided into multiple processing mods in OFDM. Each module is responsible for specific sub-tasks, such as Quadrature amplitude modulation, an inverse fast fourier transform, channel estimation, signal detection, etc. However, conventional receiver often relies on pre-set mathematical models and assumptions. It works well under ideal assumptions, but in practical applications, the channel conditions are complex and changeable (such as multi-path propagation, frequency shift, nonlinear distortion, etc.), and conventional receivers may not be able to effectively cope with these changes especially for channel estimation module.
@@
AI method can extract essential information in sparse and time-varying pilot frequencies. The complex mapping relationship between the input and the output is effectively constructed, thus AI-based receiver can learn and adapt to channel characteristics in complex environments. Each module of the conventional receiver requires accurate signal modeling and calculation, which leads to the complexity and calculation burden of the system. The AI model is able to extract features directly from the original signal and demodulate it in a data-driven manner, simplifying the complex process of conventional receivers.
@@
Figure 7: This is the difference between conventional OFDM and AI-based OFDM receivers. Using AI modules to replace conventional physical layer processing, AI’s capabilities can compensate for signal distortion.
@@
Regarding an OFDM wireless receiver, some researchers have used a fully connected deep neural network (DNN) to improve the existing modular OFDM receiver. A typical framework of AI-based receiver is given in Figure 7. We use DNN parameters $\Phi$ to construct the relationship between input $X$ and output $Y$ as $Y:f(X,\Phi)$. With the assumption of $L$ hidden layers, we denote $w_{l}$ as the parameter weights of $l-th$ layer, where $l$ ranges from 00 to $L+1$. $w_{0}$, and $w_{L+1}$ represent the weights of input layer, output layer, respectively. $b$ denotes the bias of each layer. In the training stage, the optimization direction of model is minimizing the loss between the output $Y$ and true label $Y^{true}$. Thus the network parameters $\Phi$ are updated by changing the $w_{l}$ during the process of back propagation. Hence, the AI model can encode the information of actual scenes into neural networks and optimize the performance of algorithms by adjusting the structure and parameters of neural networks. Thus, the performance of the receiver is significantly improved.
@@
In [44], a model-driven DL approach was proposed, which combines DL with expert knowledge to replace existing OFDM receivers in wireless communications. An ML-assisted physical layer receiver technique was proposed to demodulate the OFDM signals, which were subject to very high Doppler effects and corresponding distortions in the received signal [45]. To find a balance between full-size cyclic prefix (CP) and non-existent CP, the authors of [46] investigated the redundancy problem and proposed a minimum redundant OFDM receiver using DL tools. In [47], the receivers were designed based on DNN in DL, consisting of a layer of DNNs and soft decisions. The problems of channel estimation error, delay, and decoding limits between users with conventional detection methods have been solved. It can recover the symbols of all users at one time and jointly perform channel estimation and signal detection. A novel generative supervised DNN was designed in [48], which could be trained using a reasonable number of pilots. After channel estimation, the neural network-based receiver (specifically, the autoencoder) jointly learns the pre-encoder and decoder for data symbol detection. An intelligent receiver for the OFDM communication system was designed in [49] based on the DNN structure, and the intelligent receiver was realized by optimizing the DNN structure. This method can recover information at the receiving side and avoid complicated pilot operations and signal error accumulation. Moreover, a convolutional neural networks (CNN) was used to reduce the bit error rate by using the mathematical calculation function of discrete Fourier transform and the training of OFDM signal samples [50].
Further, when we replace the transmitter with AI module and train the transmitter and receiver jointly, we can realize end-to-end communication. However, in existing studies, the channel model in end-to-end communication is simple, and the existing methods are difficult to directly extend to actual deployment. Meanwhile, it is still necessary to improve the model’s generalization ability for complex wireless channels. For practical deployment, we should take into account the difficulty of solving the gradient direction propagation of online training. AI-based receivers have shown great benefits and development potential in various conditions. For actual deployment, we still need to further consider the model’s generalization ability on potential new environments, as well as its robustness in harsh environments.
@@ 2.1.3 AI-Based Beam Management
In light of the necessity for 6G networks to support elevated data rates, diminished latencies, and broader connectivity, beam management technology assumes a pivotal function in the efficient exploitation of spectrum resources, the augmentation of communication coverage, and the assurance of communication quality. The massive multi-antenna system within 6G engenders a greater quantity of beams with more constricted widths, consequently encountering novel challenges with respect to beam alignment.
@@
In the existing academic literature, the beam selection methods of multi-antenna systems can be divided into several categories, including beam selection based on beam scanning [51], beam selection based on training signals [52], beam selection based on position prediction [53], and beam selection based on hierarchical search [54]. However, the conventional beam management methods have some limitations when applied to millimeter wave massive multiple-input multiple-output (MIMO) systems [55]. The primary cause is that the real-world wireless propagation environment often has various characteristics. After a beam is narrowed, the signal transmission is more susceptible to occlusion. The increase in the number of beams significantly increases the number of beam search operations in beam alignment. It is difficult to accurately select the optimal beam instantaneously.
@@
Figure 8: This is the structure diagram of AI-based beam management. AI selects the optimal beam by looking for the complex mapping relationship between input and output, which can significantly reduce the search cost of the beam while ensuring prediction accuracy.
@@
With the application of AI, the optimal beam can be selected without scanning all the beam pairs, thus alleviating the scanning overhead existing in the conventional beam management method as shown in Figure 8. In [56], DNN was utilized for beam alignment based on the context information of the UE location. The proposed method can reduce the search time by 4x for AP selection and over 10x for beam selection. By taking advantage of the mmWave channel structure, a novel hierarchical beam alignment framework was proposed in [57] to leverage DL techniques to seek the optimal beamformer with learnable probing codebooks. This novel framework learns two tiers of probing codebooks and uses their measurements to predict the optimal beam in a coarse-to-fine search manner. A DL-based beam alignment method was proposed in [58] by jointly training the probing codebook and the beam predictor. Simulation results demonstrate that the proposed method can achieve high beam alignment accuracy while reducing the beam sweeping complexity by 10×.
@@
The above literature aims to use DL methods to improve the accuracy and robustness of beam alignment, and has achieved excellent performance, demonstrating the superiority of AI. For practical deployment, we should further consider the acquisition of high-quality tags in complex environments. Considering the training overhead, improving the generalization ability of AI models among different BS environments also becomes an important future work.
In the actual communication scenario, the massive historical beams provide enough data for training AI models, and the search for the optimal beam based on the correlation between beams constructed by AI dramatically reduces the search space of the beam. At the same time, due to the implicit structure of AI, more complex functional relationships can be explored. Compared with the conventional beam management method received signal strength (RSS), even in a low signal-to-noise ratio (SNR) environment, the optimal beam selection based on AI still has obvious advantages and robustness.
@@ 2.1.4 AI-Based Wireless Localization
The 6G network is mandated to furnish not only ultra-high-speed data transmission and ultra-low-latency communication services but also high-precision wireless location capabilities for diverse application scenarios. Ranging from the accurate positioning of vehicles in intelligent transportation systems, to the location tracking of devices in the industrial IoT, and to indoor high-precision location services, wireless location represents one of the crucial technologies for actualizing these applications. Wireless positioning of 6G refers to obtaining the location information of mobile users in a cellular environment. This location information is usually in the form of the mobile user’s geographic coordinates relative to some reference points. Wireless location is also commonly referred to as mobile location, wireless location, and geolocation [59].
Wireless positioning technology can be divided into wide and local area positioning according to the scope of services. Wide area positioning provides users with universal positioning services. Local area positioning is mainly used in indoor scenarios. Wide area positioning primarily includes mobile communication network positioning and global satellite navigation systems, e.g., the global positioning system (GPS). The GPS is a kind of high-precision radio navigation positioning system based on artificial Earth satellites, which can provide accurate geographical position, vehicle speed and accurate time information anywhere on the Earth and in near-earth space. Local wireless positioning mainly includes Wi-Fi, Bluetooth, UWB, etc. Its positioning technology can be divided into four types: ranging positioning based on RSS, fingerprint positioning based on signal characteristics, positioning based on signal transmission time and angle of arrival measurement.
Many applications need high-precision positioning, such as industrial automated guided vehicle and asset tracking, especially indoor precision positioning, but GPS cannot be used indoors. Despite having a specific positioning reference signal, LTE cannot meet the requirements for high-precision positioning due to its relatively low positioning accuracy and the large distance between BS (about 100 meters even with a 20 MHz bandwidth). Bluetooth, Wi-Fi and other wireless location-based technologies have high deployment costs and are difficult to become a universal positioning technology [60].
The application of AI technology has been introduced to achieve high-precision positioning in the 5G era. One approach is to utilize the random forest technique [61], which can create a classification model to split a vast area into many small grids and then predict the grid where the user is located. It can also make a regression model to predict the user’s position coordinates. In another method, multi-layer perceptron has a similar application to random forest and can establish regression and classification models to locate targets [62]. AI algorithms can also be trained to recognize the fingerprint of a wireless signal in a specific location and process and fuse data from different sensors to achieve high-precision positioning.
[…]

### v4

@@ 2 AI for Network
The most fundamental goal of integrating wireless communication and AI lies in enhancing the performance, and efficiency of the network and the service experience of users by virtue of AI, which is known as AI4NET [39]. At this stage, the focus is on how to utilize AI algorithms to optimize communication performance and network functions. For example, AI can optimize the signal modulation and demodulation process, making signal transmission more accurate and efficient; with the help of AI, network resources can be intelligently allocated to achieve load balancing; and an automated O&M management mode can be constructed to improve O&M efficiency and reduce costs. This introduction of AI is not anticipated to have a significant impact on the original network architecture. Instead, it is to improve the communication problems in specific recognition by training AI algorithm models. This process is analogous to precisely implanting intelligent patches in the network. It can not only maintain the stability of the network architecture but also gradually enhance the intelligence level of the network.
@@
The concept of AI4NET is presented in Figure 5, where the wireless AI possesses capabilities such as feature extraction, prediction, adaptation, optimization, real-time processing, correlation, and scene clustering. These capabilities form the foundation for AI-enabled wireless networks and can be directly deployed on the BS side and the CN to empower wireless communication. AI optimizes conventional methods to enhance network performance on various terminal devices. On the network side, AI takes advantage of its capabilities to improve the quality of end-to-end services, optimize the functions of network elements, conduct automated management of the upper-layer network, and boost network O&M efficiency. The advent of AI has also given rise to a series of new services aimed at improving and serving users in a better manner.
AI4NET has initiated relevant research and applications in 5G. In 6G, with the more mature AI technologies represented by DL and the emergence of new infrastructures integrating connectivity and computing power, the relevant applications will become more abundant and sophisticated and may further evolve in depth. This section elaborates on some cases from two aspects, namely air interface performance enhancement and network O&M improvement, to illustrate the concept of AI4NET.
@@ 2.1 Air Interface Performance Enhancement
In this subsection, we show how AI technology can improve transmission in channel state information (CSI) feedback, orthogonal frequency division multiplexing (OFDM) receivers, beam management, and wireless localization.
@@ 2.1.1 AI-Based CSI Feedback Algorithm
Within the 6G framework, the feedback overhead is expected to experience a sharp increase along with a significant rise in the number of antennas. Consequently, ensuring the accuracy of channel reconstruction and minimizing the CSI feedback cost is one of the bottleneck problems that need to be surmounted. The utilization of AI technology in CSI feedback can profoundly dissect these elaborate channel characteristics. Through training with a substantial quantity of labeled samples, the AI model can acquire the characteristic patterns of the channel within diverse fading environments, interference scenarios, and multipath circumstances. In contrast to conventional CSI estimation approaches relying on statistical theory [40] or simple linear models [41], AI-based methods exhibit stronger generalization competencies and can estimate CSI with greater precision in complex real-world communication settings.
The massive CSI data and the inherent random characteristics of CSI make AI potentially useful for designing a new CSI feedback mechanism. For instance, the nonlinear characteristics of DL can be utilized to efficiently extract the features of CSI, and the original CSI data can be transformed into a more compact feature representation, thereby reducing the communication overhead in the CSI feedback process. An example of its implementation is the autoencoder architecture, which employs neural networks to extract and compress the features of CSI. The authors of [42] were the first to apply DL to CSI feedback. The proposed CsiNet scheme utilizes convolution to extract channel features, and compress and reconstruct channels, and it is superior to the CSI feedback scheme based on compressed sensing in terms of feedback accuracy and computational complexity. The encoder of CsiNet compiles the high-dimensional CSI into code words, while the decoder is responsible for decoding the codewords back into the original CSI. The internal structures of encoders and decoders of different algorithms vary, resulting in differences in channel reconstruction performance. As shown in Figure 6, Huawei demonstrated the huge advantages of AI-based CSI feedback over conventional codebook approaches through system-level simulations. AI with additional information can effectively improve the performance on the square of the generalized cosine similarity (SGCS) and throughput [43].
@@
Figure 6: The performance of R16 Type II codebook, AI-SF, and AI-SFT on SGCS and throughput gain. AI-SF means the AI/ML-based spatial-frequency compression method, which transformer is used as the backbone of both the encoder and decoder. AI-SFT is AI-SF with additional past CSI.
The studies in [44] increased the size of the convolutional kernel based on CsiNet to improve the perceptual field of view of the convolutional layer, which is conducive to extracting the sparsity of the channel and enhancing the accuracy of channel reconstruction. Although a larger perceptive field of view could effectively extract the sparsity of CSI, a smaller convolution kernel could extract finer features from CSI. A DL-based CSI compression and quantization method was developed in [45] considering high compression ratios. A multi-rate CSI compression framework was designed in [46] to improve the generalization of the model in the field of transfer learning. In [47], a non-local neural network was introduced based on the CsiNet network to capture a wide range of dependencies, and the accuracy of channel recovery was improved compared to CsiNet. The CRNet was proposed in [48], which used convolution kernels of different sizes for channel feature extraction and recovery, reducing the computation and improving the accuracy of channel reconstruction. In order to better deploy AI-based CSI feedback models in real-world wireless environments, model monitoring, model updating and AI-based signaling process management are hot issues that require our attention in future studies.
@@ 2.1.2 AI-Based OFDM Receiver
AI method can extract essential information in sparse and time-varying pilot frequencies. The complex mapping relationship between the input and the output is effectively constructed, thus AI-based receiver can learn and adapt to channel characteristics in complex environments. Each module of the conventional receiver requires accurate signal modeling and calculation, leading to the complexity and calculation burden of the system. AI has the potential to extract features directly from the original signal and demodulate it in a data-driven manner, simplifying the complex process of conventional receivers.
@@
Figure 7: The AI-based OFDM receivers. Using AI modules to replace conventional physical layer processing, AI’s capabilities can compensate for signal distortion.
@@
Regarding wireless OFDM receivers, some researchers have used a fully connected deep neural network (DNN) to improve the existing modular OFDM receiver. A typical framework of an AI-based receiver is given in Figure 7. We use DNN parameters $\Phi$ to construct the relationship between the input $X$ and output $Y$ as $Y:f(X,\Phi)$. With the assumption of $L$ hidden layers, we denote $w_{l}$ as the parameter weights of the $l$-th layer, where $l$ ranges from 00 to $L+1$. $w_{0}$ and $w_{L+1}$ represent the weights of the input layer and the output layer, respectively. $b$ denotes the bias of each layer. In the training stage, the optimization direction of the model minimizes the loss between the output $Y$ and true label $Y^{true}$. Thus, the network parameters $\Phi$ are updated by changing $w_{l}$ during the process of back propagation. Hence, the AI model can encode the information of actual scenes into neural networks and optimize the performance of algorithms by adjusting the structure and parameters of neural networks.
@@
In [49], a model-driven DL approach was proposed, which combines DL with expert knowledge to replace existing OFDM receivers in wireless communications. An ML-assisted physical layer receiver technique was proposed to demodulate the OFDM signals, subject to very high Doppler effects and corresponding distortions in the received signal [50]. To strike a balance between full-size cyclic prefix (CP) and non-existent CP, the authors of [51] investigated the redundancy problem and proposed a minimum redundant OFDM receiver using DL tools. In [52], the receivers were designed based on DNN, consisting of a layer of DNNs and soft decisions. The problems of channel estimation error, delay, and decoding limits between users with conventional detection methods were solved. The symbols of all users were recovered at one go to jointly perform channel estimation and signal detection. A novel generative supervised DNN was designed in [53], which could be trained using a reasonable number of pilots. After channel estimation, the neural network-based receiver jointly learns the pre-encoder and decoder for data symbol detection. An intelligent receiver for OFDM communication systems was designed in [54] based on the DNN structure, and realized by optimizing the DNN structure. This method can recover information on the receiving side and avoid complicated pilot operations and signal error accumulation. Moreover, a convolutional neural network (CNN) was used to reduce the bit error rate using the mathematical calculation function of discrete Fourier transform and the training of OFDM signal samples [55].
@@ 2.1.3 AI-Based Beam Management
In the existing literature, the beam selection methods of multi-antenna systems can be divided into beam selection based on beam scanning [56], beam selection based on position prediction [57], and beam selection based on hierarchical search [58]. Conventional beam management methods have some limitations when applied to millimeter wave massive multiple-input multiple-output (MIMO) systems [59]. The primary cause is that the real-world wireless propagation environment often has various characteristics. After a beam is narrowed, the signal transmission is more susceptible to occlusion. The increase in the number of beams significantly increases the number of beam search operations in beam alignment. It is difficult to accurately select the optimal beam instantaneously.
@@
Figure 8: The structure diagram of AI-based beam management. AI selects the optimal beam by looking for the complex mapping relationship between input and output, which can significantly reduce the search cost of the beam while ensuring prediction accuracy.
@@
With the application of AI, the optimal beam can be selected without scanning all beam pairs, thus alleviating the scanning overhead existing in the conventional beam management method, as shown in Figure 8. In [60], a DNN was utilized for beam alignment based on the contextual information of the UE location. This method can reduce the search time by four times for AP selection and over tenfold for beam selection. By taking advantage of the mmWave channel structure, a novel hierarchical beam alignment framework was proposed in [61] to leverage DL techniques to seek the optimal beamformer with learnable probing codebooks. This novel framework learns two tiers of probing codebooks and uses their measurements to predict the optimal beam in a coarse-to-fine search manner. A DL-based beam alignment method was proposed in [62] by jointly training the probing codebook and the beam predictor. Simulation results demonstrate that the method can achieve high beam alignment accuracy while reducing the beam sweeping complexity by ten times.
@@
The above literature aims to use DL methods to improve the accuracy and robustness of beam alignment, and has demonstrated the superiority of AI. For practical deployment, we should further consider the acquisition of high-quality tags in complex environments. Considering the training overhead, improving the generalization ability of AI models among different BS environments also becomes an important future work.
@@ 2.1.4 AI-Based Wireless Localization
The 6G network is mandated to furnish high-precision wireless location capabilities for diverse application scenarios. This location information is usually in the form of the mobile user’s geographic coordinates relative to some reference points [63]. Many applications need high-precision positioning, such as industrial automated guided vehicle and asset tracking, especially indoor precision positioning, but the global positioning system (GPS) cannot be used indoors. Despite having a specific positioning reference signal, LTE cannot meet the requirements for high-precision positioning due to its relatively low positioning accuracy and the large distance between BS (about 100 meters even with a 20 MHz bandwidth). Bluetooth, Wi-Fi, and other wireless location-based technologies can have high deployment costs and are difficult to become a universal positioning technology [64].
The application of AI technology has been introduced to achieve high-precision positioning in the 5G era. One approach is to utilize the random forest technique [65], which can create a classification model to split a vast area into many small grids and then predict the grid where the user is located. It can also make a regression model to predict the user’s position coordinates. In another method, multi-layer perceptron has a similar application to random forest and can establish regression and classification models to locate targets [66]. AI algorithms can also be trained to recognize the fingerprint of a wireless signal in a specific location and fuse data from different sensors to achieve high-precision positioning.
With the widespread use of massive MIMO technology, AI has shown more significant advantages in high-precision positioning. Using the location in the line-of-sight (LoS) and non-line-of-sight (NLoS) channel co-existence scenario based on the neural network as an example, ML and a significant volume of channel data can help effectively map the relationship between channel response [67] and position coordinates, resolving the location problem in complex environments and increasing location accuracy. Very high positioning accuracy is achieved without significant additional cost, even about 20 mm accuracy under indoor LoS and NLoS conditions. Due to the diversity of data modes in complex environments, the further development and practical application of AI-based localization requires us to consider the robustness and reliability of the AI model to enable mobility management.